OpenStudy (anonymous):

Water is leaving a hose at 6.8m/s . If the target is 2m away horizontally, what angle should the water leave the base.

OpenStudy (binary3i):

hint let the angle be \[\theta\] now make it in to X and Y components. acceleration is in negative Y direction which is -g. so use equation V=U+at. now try.......

OpenStudy (souvik):

let the angle be \[theta\] and the time taken to hit the target is T X component of the velocity is \[v \cos \theta \] then the distance between pipe and the target is \[v \cos \theta T \] now....\[v \cos \theta T =2\]...............................(1) Y component of the velocity is v.sin(theta) and the acceleration is -g as it is the opposite direction of velocity... we use this equation.... \[h _{y}=u _{y}t-1/2g t ^{2}\] here \[u _{y}=u \sin \theta \] and after time T water will come back at it initial point..so h=0 after time T \[0=usin \theta T-1/2g T ^{2}\] or \[T=2usin \theta /g\].................................(2) eliminate T from (1) and (2)...you will get the answer.. \[\theta=12.5^{0}\]

OpenStudy (binary3i):

@souvik you shoul not tell the answers directly with out letting them to attempt solving it.